40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Jan 20 20:20:36 2020, gautam, Update, PSL, PMC servo checkout PMCsweep.pdfOLTFmeas.pdf
    Reply  Wed Jan 22 11:22:39 2020, gautam, Update, PSL, PMC modulation depth measurement modDepth.pdf
       Reply  Wed Jan 22 20:12:36 2020, gautam, Update, PSL, PMC demodulator electrical characterization demodChar.pdfmixerChar.pdf
          Reply  Thu Jan 23 14:37:05 2020, gautam, Update, PSL, PMC VGA chip damaged? VGAchar.pdf
             Reply  Thu Jan 23 16:37:14 2020, rana, Update, PSL, PMC VGA chip damaged? 
             Reply  Thu Jan 23 22:10:01 2020, gautam, Update, PSL, PMC servo pulled out 
                Reply  Fri Jan 24 15:42:08 2020, gautam, Update, PSL, PMC servo restored 
                   Reply  Sun Jan 26 13:47:00 2020, gautam, Update, PSL, PMC servo characterization elecTFs.pdfVGAchar_postFix.pdfVGAlinearity_postFix.pdfnewOLTFs.pdf
                      Reply  Tue Jan 28 14:33:24 2020, gautam, Update, PSL, Inferred free-running frequency noise inLoopNoise_IMClocked.pdffreqNoiseComparison.pdf
          Reply  Thu Jan 23 18:52:31 2020, gautam, Update, PSL, PMC RFPD characterization PDresp.pdf
Message ID: 15152     Entry time: Fri Jan 24 15:42:08 2020     In reply to: 15149     Reply to this: 15156
Author: gautam 
Type: Update 
Category: PSL 
Subject: PMC servo restored 

The PMC servo was re-installed at ~345pm. HV supplies were re-energized to their nominal values. I will update the results of the investigation shortly. The new nominal PMC servo gain is +9dB.

Quote:

While I have the board out, I'll try and do a thorough investigation of TFs and noise of the various stages. There is no light into the IFO until this is done.

I pulled the board out at 345pm after dialling down all the HV supplies in 1X1. I will reinstall it after running some tests.

ELOG V3.1.3-